OpenStudy (anonymous):

A line passes through (3, –2) and (6, 2). Write an equation for the line in point-slope form. Rewrite the equation in standard form using integers.

OpenStudy (anonymous):

y – 2 = mc013-4.jpg(x – 3); –4x + 3y = 18 y + 2 = mc013-2.jpg(x + 3); –4x + 3y = –18 y – 3 = mc013-3.jpg(x + 2); –4x + 3y = 17 y + 2 = mc013-1.jpg(x – 3); –4x + 3y = –18